OpenStudy (anonymous):

f(x)=xe^(2x) find [f(a^2)]^2

OpenStudy (dape):

\[ f(a^2) = a^2*e^{2a^2} \] \[ (f(a^2))^2 = \left(a^2*e^{2a^2}\right)^2 = (a^2)^2*(e^{2a^2})^2 = a^{2*2}*e^{2*2a^2} = a^4e^{4a^2} \]